OpenStudy (anonymous):

A cylindrical vase has a diameter of 6 inches. At the bottom of the vase, there are 9 marbles, each of diameter 2 inches. The vase is filled with water up to a height of 12 inches. What is the volume of water in the vase? 96π in3 27π in3 65π in3 105π in3

OpenStudy (btaylor):

calculate the volume of the marbles (V = 4/3 π r^3), then multiply that by nine. calculate the volume of the vase (V = π r^2 h), then subtract the marbles' volume.

OpenStudy (anonymous):

π(3^2)(12)??

OpenStudy (btaylor):

but aren't there only 9 marbles? You put 12.

OpenStudy (anonymous):

12=h

OpenStudy (campbell_st):

marble volume \[V = 9 \times \frac{4}{3} \pi \times1^3\] \[V = 12 \pi\] volume of the vase \[V = \pi \times 3^2 \times 12\] \[V = 108 \pi\] Volume of water \[V = 108 \pi - 12 \pi = 96 \pi\]

OpenStudy (anonymous):

thank you both:)

OpenStudy (anonymous):

calculate volume of each marble and then total volume occupied by the marbles. now, actual volume of water = volume of water column (i.e. up to the ht of water level) - total volume occupied by the marbles

OpenStudy (anonymous):

thanks @annas

OpenStudy (anonymous):

my pleasure :)

OpenStudy (shaw74):

is the answer 96 pie